OpenStudy (anonymous):

TASK 1 solve using diffident methods. explained which method you found to be more efficient a. 3x - 9y = 3 6x - 3y = -24 b. 7x - 3y = 20 5x + 3y = 16 c. y - 1/2x = 6 2x + 6y = 19

OpenStudy (anonymous):

a. method 1 multiply top by 2 then subtract 2*(3x-9y) = 2*3 - 6x -3y =-24 -15y = 30 y = -2 6x + 6 = -24 x = -5 method 2 divide top by -3 then subtract method 3 multiply bottom by -3 then add I'm sure there are more. Do you see how I did the first? I will help you do the rest if I am on and you don't understand.